# RoundaboutHD: High-Resolution Real-World Urban Environment Benchmark for Multi-Camera Vehicle Tracking

### v1

@@ 4.2 Single Camera Tracking
In the general MTMCT task, the performance of MTMCT largely depends on the performance of SCT, as errors in SCT can negatively affect downstream association across cameras. As a result, many researchers have focused on improving SCT to enhance the overall effectiveness of MTMCT systems. In our RoundaboutHD dataset, we provide some extra challenging for SCT. Due to the nature of roundabout, vehicles follow non-linear, curved trajectories rather than simple straight-line motion. In addition, a central statue located at the center of the roundabout (visible in Fig. 2) introduces frequent occlusions. These factors collectively increase the difficulty of tracking consistent vehicle identities over time within each camera view.
To evaluate SCT performance on RoundaboutHD, we follow the widely used tracking-by-detection paradigm provided by BOXMOT [4]. BOXMOT is an open-source pluggable SCT framework that enables easy and straightforward deployment of multiple different SOTA SCT algorithms with minimal configuration change. For consistency across evaluations, detection results from the YOLOv12x model are used as input for all trackers. This ensures that the evaluation focus on the performance differences of tracking algorithms themselves. Additionally, to avoid distractions from static vehicles parked in visible parking areas, trajectories corresponding to stationary objects are filtered out prior to evaluation. Detailed results from different trackers are listed from Tab. 6 to Tab. 10, and a performance summary is shown in Tab. 11.
@@
Among all evaluated algorithms, ByteTrack achieves the highest average performance in terms of both IDF1 and MOTA. BoostTrack reports the lowest number of ID switches; however, this might because of its limited tracking performance, as indicated by its lower IDR score, which suggests fewer trajectories are counted in the result. For the comparison among different scene, trackers generally perform better in Cam02 and Cam04 compared to Cam01 and Cam03. This disparity is largely due to the presence of a large statue located at the center of the roundabout in Cam01 and Cam03, which blocks the view and introduce frequent occlusion during the tracking progress.
@@ 4.3 Image-Based ReID
To evaluate the image-based vehicle ReID subset, we adopt the widely-used FastReID toolbox [14], which provides a straightforward framework for training and evaluating various SOTA algorithms. We select three popular baseline models commonly used for ReID task, BoT [24], AGW [37] and SBS (a FastReId improved baseline based on BoT) [14] for benchmarking the RoundaboutHD image-based ReID subset. To ensure a fair comparison, all three baselines utilize the same ResNet50-IBN backbone, pre-trained on ImageNet [8]. Each baseline model is trained on our training dataset using the combination of CrossEntropyLoss and TripletLoss with FastReID’s default training parameters for 60 epochs, and the best-performing model is selected for the final evaluation. All training is conducted on an NVIDIA GeForce RTX 4070 Ti GPU. The results are shown in Tab. 12 and Tab. 13. Two quantitive examples for image-based ReID are visualized by Fig. 7.
Figure 7: Rank-50 Retrieval Example for Image-Based Vehicle ReID using the SBS [14] Model Trained on the RoundaboutHD Subset. The blue bounding box marks the query image, green bounding boxes indicate correctly retrieved matches (true positives), and red bounding boxes denote incorrect matches (false negatives).
@@
Table 13: Pre-trained Vehicle ReID performance. This table shows the performance of models trained on our custom dataset, all values are in percentage.
As shown in Tab. 12, all three models initially achieve low mAP when using only ImageNet-pretrained weights, highlighting the domain gap between generic image datasets and real-world vehicle ReID tasks. However, fine-tuning on the RoundaboutHD dataset leads to substantial performance gains across all models. As shown in Fig. 7, the trained model achieves near-perfect accuracy in matching vehicle instances captured from multiple viewpoints. Notably, most false negatives are caused by the limited number of query images per identity rather than inherent model deficiencies. This significant improvement can be largely attributed to the roundabout scenario, which naturally captures vehicles from diverse angles—even within the same camera—thus enriching the visual appearance space and facilitating more robust feature learning for image-based ReID tasks.
@@ 4.4 Multi-Camera Tracking
Multi-camera tracking (MCT) is the task for associating vehicle identities across different camera views. In most of studies, MCT builds upon the outputs of SCT, associating identities by comparing distances between deep embedding features. Some approaches also integrate camera topology and spatiotemporal reasoning to further constrain the search space, enabling more accurate reconstruction of complete cross-camera trajectories.
Following the release of CityFlow and the CVPR AI City Challenge workshop, there are several research focused on the MCVT problem and achieved good results. Notable examples include ELECTRICITY [27], the winning method of the 2020 AI City Challenge, Liu et al. [19], winner in 2021, and Yang et al. [36], winner in 2022. In our benchmark evaluation, we choose ELECTRICITY as the baseline method due to its general and reproducible pipeline. Unlike Liu and Yang’s approaches, which rely on manually annotated spatial-temporal zone-based information between cameras that are only available for the CityFlow dataset, ELECTRICITY can be more broadly applied to other datasets without requiring prior human annotations. During the evaluation on RoundaboutHD, we filter out all static vehicle trajectories to remove the influence of parked vehicles visible in the scene. For the cross-camera matching stage, we fine-tune the threshold distance to 12 and set a hard removal distance of 80. The performance of ELECTRICITY across multiple datasets is summarized in Tab. 14.
@@
This result provides a baseline of how SOTA algorithm performances in different MCVT dataset, with IDF1 score serving as the primary metric. Moreover, it also proves the increased challenge posed by RoundaboutHD compared to existing dataset, which is largely attributed to nonlinear vehicle trajectories and frequent occlusions inherent in roundabout environments.
@@ 5 Conclusion
This paper presents RoundaboutHD, a comprehensive, high-resolution real-world dataset designed to advance research in MCVT. Unlike previous open-source datasets, RoundaboutHD addresses the lack of high-quality, practically representative data in this field. It introduces additional challenges such as nonlinear vehicle motion and frequent occlusions, which arise naturally from the roundabout scenario.
The dataset comprises video footage from four non-overlapping cameras, each offering 10 minutes of fully annotated 4K, 15 fps video. In total, it includes 549,909 annotated bounding boxes across 512 unique vehicle identities. In addition to the full MCVT dataset, we provide dedicated subsets for object detection, single-camera tracking, and image-based vehicle ReID. A unified evaluation framework is also released, along with baseline results using several SOTA algorithms across all tasks. Experimental results prove that RoundaboutHD presents realistic and substantial challenges, making it a valuable benchmark for the development and evaluation of future MCVT methods.
While RoundaboutHD provides high-quality and richly annotated data, its focus on a single scene may limit scenario diversity, which could affect model generalizability in more varied environments. To address this, future extensions will aim to incorporate a wider variety of conditions, including different lighting, weather, and traffic patterns. We hope that the release of RoundaboutHD will inspire further advancements in the MTMCT research community.

### v2

@@ 4.2 Single Camera Tracking
In the general MTMCT task, the performance of MTMCT largely depends on the performance of SCT, as errors in SCT can negatively affect downstream across cameras association. As a result, many researchers have focused on improving the SCT performance to enhance overall effectiveness of MTMCT systems. In our RoundaboutHD dataset, we provide some extra challenging for SCT. Due to the nature of roundabout, vehicles follow non-linear, curved trajectories rather than simple straight-line motion. In addition, a central statue located at the center of the roundabout (visible in Fig. 2) introduces frequent occlusions. These conditions make it more difficult to maintain consistent identity tracking over time within a single camera view.
To benchmark SCT performance on each camera of the RoundaboutHD dataset, we follow the widely used tracking-by-detection paradigm provided by BOXMOT [4]. BOXMOT is an open-source pluggable SCT framework that enables easy and straightforward deployment of multiple different SOTA SCT algorithms. In our evaluation, all trackers use the same set of raw detection results generated by the YOLOv12x model, ensuring that performance differences can be attributed solely to the tracking algorithms themselves. Additionally, to make a more representative evaluation, trajectories corresponding to parked vehicles in visible parking areas are filtered out prior to evaluation. Detailed results from different trackers are listed from Tab. 6 to Tab. 10, and a performance summary is shown in Tab. 11.
@@
Among all tested algorithms, ByteTrack achieves the highest average performance in terms of both IDF1 and MOTA, which are the most representative metrics for assessing overall tracking accuracy and robustness. BoostTrack reports the lowest number of ID switches; however, this might because of its limited tracking performance, as proved by its lower IDF1 and IDR score, which suggests fewer trajectories are successfully tracked. For comparison among different scenes, trackers generally perform better in Cam02 and Cam04 compared to Cam01 and Cam03. This difference is largely due to the presence of a large statue located at the center of the roundabout in Cam01 and Cam03, which blocks the view and introduces frequent occlusion during the tracking progress.
@@ 4.3 Image-Based ReID
To benchmark the RoundaboutHD image-based ReID subset, we evaluate three commonly used baseline models: BoT [24], AGW [37], and SBS (a stronger baseline built upon BoT) [14], all implemented using the FastReID framework [14] with a shared ResNet50-IBN backbone. Each model is evaluated under two settings: (1) directly using ImageNet-pretrained weights [8], and (2) after fine-tuning on our custom RoundaboutHD training set. For the fine-tuning process, we follow FastReID’s default training configuration, using a combination of CrossEntropyLoss and TripletLoss. Models are trained for 60 epochs, and the best-performing checkpoint based on validation accuracy is selected for final evaluation. All training is conducted on an NVIDIA GeForce RTX 4070 Ti GPU. The quantitative results for both settings are presented in Tab. 12 and Tab. 13, while Fig. 7 illustrates two visual examples from the ReID benchmark.
Figure 7: Rank-50 Retrieval Example for Image-Based Vehicle ReID using the SBS [14] Model Trained on the RoundaboutHD Subset. The blue bounding box on the left marks the query image, green bounding boxes indicate correctly retrieved matches (true positives), and red bounding boxes denote incorrect matches (false negatives).
@@
Table 13: Vehicle ReID performance. This table shows the performance of models re-trained on our custom dataset, all values are in percentage.
According to the benchmarking results, Tab. 12 reports the performance of the three baseline models using only ImageNet-pretrained weights, while Tab. 13 presents the results after re-training on the RoundaboutHD ReID subset. Comparing Tab. 12 and Tab. 13, the re-trained model shows an obvious advantage in terms of our selected metrics. This result highlights the domain gap between generic image datasets and real-world vehicle ReID tasks. Meanwhile, the huge performance increment also proves the effective and comprehensive of the RoundaboutHD ReID subset, which provides vehicle images shot from diverse angles and facilitates more robust feature learning for image-based ReID tasks. However, even with strong overall performance on re-trained model which reaches nearly perfect scores in mAP, Rank-5, and mINP, challenges still remain at larger retrieval depths. As shown in Fig. 7, the top example demonstrates excellent retrieval of a blue Mack truck, with 100% AP on the rank-50 result. In contrast, the bottom example shows a limited performance for a white Tesla Model Y on rank 19th to rank 50th. This suggests the model finds rare or visually distinctive vehicles more easily than common ones. This is also consistent with the vehicle distribution in Fig. 4, where only 7.2% of vehicles are blue and 5.4% are HGVs, compared to 27.6% white vehicles and 45.6% SUVs, which are more common and harder to distinguish.
@@ 4.4 Multi-Camera Tracking
Multi-camera tracking (MCT) is the task for associating vehicle identities across different camera views. In most of studies, MCT builds upon the outputs of SCT, associating identities by comparing distances between deep embedding features. Some approaches further incorporate camera topology and spatiotemporal reasoning—using inter-camera distances to estimate transition probabilities and narrow down the searching space for more accurate cross-camera association.
Following the release of CityFlow and the CVPR AI City Challenge workshop, there has been significant several efforts focusing on the MCVT problem and achieving good results. Notable examples include ELECTRICITY [27], the winning method of the 2020 AI City Challenge, Liu et al. [19], winner in 2021, and Yang et al. [36], winner in 2022. In our benchmark evaluation, we choose ELECTRICITY as the baseline method due to its general and reproducible pipeline. Unlike Liu and Yang’s approaches, which rely on manually annotated spatial-temporal zone-based information between cameras that are only available for the CityFlow dataset, ELECTRICITY can be more broadly applied to other datasets without requiring prior human annotations. During the evaluation on RoundaboutHD, we filter out all static vehicle trajectories to remove the influence of parked vehicles visible in the scene. For the cross-camera matching stage, we fine-tune the threshold distance to 12 and set a hard removal distance of 80. The performance of ELECTRICITY across multiple datasets is summarized in Tab. 14.
@@
This result provides a baseline of how SOTA algorithm performances in different MCVT datasets, with IDF1 score serving as the primary metric. Moreover, it also reinforces the increased challenge posed by RoundaboutHD compared to existing dataset, which is largely attributed to nonlinear vehicle trajectories and frequent occlusions inherent in roundabout environments.
@@ 5 Conclusion
This paper presents RoundaboutHD, a comprehensive, high-resolution real-world dataset designed to advance research in MCVT. Unlike previous open-source datasets, RoundaboutHD addresses the lack of high-quality dataset that closely reflect real-world conditions, and it also introduces additional challenges such as nonlinear vehicle motion and frequent occlusions, which arise naturally from the roundabout scenario.
The dataset comprises video footage from four non-overlapping FOV cameras, each offering 10 minutes of fully annotated 4K, 15 fps video. In total, it includes 549,909 annotated bounding boxes across 512 unique vehicle identities. In addition to the full MCVT dataset, we also provide subsets for object detection, single-camera tracking, and image-based vehicle ReID. The dataset further includes camera modeling/geometry and vehicle-level contextual information(vehicle type, color, make and model), which enriches its applicability to multi-modal and multi-task scenarios. To support standardized evaluation, we provide a unified evaluation framework along with baseline results from several SOTA algorithms across all tasks. These experimental results prove that RoundaboutHD presents realistic and substantial challenges, making it a valuable dataset for the evaluation of future MCVT methods.
Despite its strengths, the focus on a single scene may limit the dataset scenario diversity, which may affect model generalization. To address this, future extensions will aim to incorporate a boarder range of conditions, including different lighting, weather, and traffic patterns. We hope that the release of RoundaboutHD will inspire further advancements in the MTMCT research community.